OpenStudy (anonymous):

A clock gains 2 minutes per day and another loses 3 minutes per day. Both the clocks started on a correct time at a given instant. They will again both show the correct time for an nstant within a. a week b. a month c. six months d. one year

OpenStudy (lgbasallote):

i see why the owner just doesnt replace the batteries so it will run right =_= why must they make it complicated

OpenStudy (anonymous):

Math is complicated.

OpenStudy (lgbasallote):

dont have to be :/ them mathematicians jjust love to see the world burn

OpenStudy (anonymous):

PLS help I have an exam tomorrow.

OpenStudy (anonymous):

Write an expression for times when the first clock will be correct. Then write an expression for times when the second clock will be correct. Set them equal to 0. Solve.

OpenStudy (anonymous):

I'm confused, how do you do that?

OpenStudy (anonymous):

The first clock gains 2 minutes per day. After 30 days, it will be 60 minutes ahead, or a full hour. To be correct again, it would need to gain 12 hours. How many days would it take to do that?

OpenStudy (anonymous):

...

OpenStudy (anonymous):

(12*60)/2 = 6 * 60 = 360 days.

OpenStudy (anonymous):

It gains 2 minutes per day. 6 days would gain it 12 minutes.

OpenStudy (anonymous):

Okay yes. So that means that every 360 days, the first clock is correct. Do the same thing for the second clock.

OpenStudy (unklerhaukus):

a stopped clock is more accurate (accurate twice a day) than a slow clock

OpenStudy (anonymous):

The second clock loses 3 minutes per day. So it take 20 days to lose an hour, and 240 days to lose 12 hours, which would make it correct again. The first clock is correct every 360 days, and the second clock is correct every 240 days.